OpenStudy (anonymous):

Slope Intercept Form of the equation of the line? y - 4 = 3(x - 7)

jimthompson5910 (jim_thompson5910):

y - 4 = 3(x - 7) y - 4 = 3x - 21 y = 3x - 21+4 y = 3x - 17